OpenStudy (anonymous):

will award medal. Indicate in standard form the equation of the line through the given points. P(0, -4), Q(5, 1) ___=______

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Hey @Bookworm101m are you also bookworm00981?

OpenStudy (anonymous):

um I don't think so

OpenStudy (anonymous):

Anyways use y- y,= m(x - x,) so plug in the numbers

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

sorry this is kinda late my stuff froze so u do y-0=1(x+4). M is basically the slope. then distribute

OpenStudy (anonymous):

whats the equation?

OpenStudy (anonymous):

y-0=1(x+4) distribute the 1then put all the variables to one side

OpenStudy (anonymous):

standard form is Ax+By=C

OpenStudy (anonymous):

im still unsure what my answer should be

OpenStudy (anonymous):

y-0=1(x+4) y-0=x+4 y-x=4